OpenStudy (anonymous):

One factor of 12t^2 − 60t + 48 is t − 1. Find the other factors, then write the polynomial in factored form.

OpenStudy (anonymous):

@thewonderfuladele, I think I kinda get this one...maybe lol

OpenStudy (anonymous):

Ok, so one factor is already there so that means, GCF would be 12, right?

OpenStudy (ranga):

Factor out 12 first. Then use the AC method to factor.

OpenStudy (anonymous):

ok, hold on.

OpenStudy (anonymous):

wait, t, not x

OpenStudy (ranga):

12t^2 − 60t + 48 = 12(t^2 - 5t + 4) = 12(t -1)(t ? )

OpenStudy (anonymous):

12(t-1)(t+4)?

OpenStudy (ranga):

t - 4

OpenStudy (anonymous):

omg, I was gonna say -4, but I always get positives & negatives mixed up

OpenStudy (ranga):

You can also solve this by synthetic division: Divide 12 - 60 + 48 by +1 (since t-1 is a factor)

OpenStudy (ranga):

1 | 12 - 60 + 48 | +12 - 48 |__________________________ 12 -48 0 So (12t - 48) is the other factor. Factor out 12: 12(t - 4) So the factored form is: 12(t-4)(t-1)